OpenStudy (anonymous):

What is an equation of the line in slope-intercept form? m= 3 and the y-intercept is (0, 4) (1 point) y= 4x+ 3 y= 3x+ 4 y= 4x3 y= 3x+ 4

OpenStudy (anonymous):

@babyblue123 @iambatman @One098 @poopsiedoodle @sammixboo

OpenStudy (anonymous):

Thanks for viewing my question!!! :D

OpenStudy (anonymous):

Do you know how solve this? \[y-4=3(x-0)\] If so, this will lead you to your answer.

OpenStudy (anonymous):

Not really, could you walk me through it?

OpenStudy (anonymous):

Original Equation: \[y-4=3(x-0)\] Multiply 3 to the parenthesis, and you'll receive: \[y-4=3x-0\] Divide 3 to both sides of the equation, you'll receive: \[=-1.33(Repeating)\]

OpenStudy (anonymous):

-1.33 is the answer? (thanks for showing the steps by the way)

OpenStudy (anonymous):

@vzfreakz ?

OpenStudy (anonymous):

@vzfreakz please help me finish???

OpenStudy (anonymous):

I guess so.

OpenStudy (anonymous):

okay so whats the next step? @vzfreakz

OpenStudy (anonymous):

That's the step I've shown you above.

OpenStudy (anonymous):

The overall answer would be, 3x+4. Why do you have two same answer choices? B and D?

OpenStudy (anonymous):

Oh sorry, B is actually -3x+4. So it's D then, right? @vzfreakz